if you have @ angles that are supplemantary, and one is 5 degree more than 6 times the other. what is the larger angle?
Answer by mananth(16949)

let one angle be x
the other angle will be 6x+5
x+6x+5=180
7x=175
x=25
6x+5 = 150+5=155
155 is the bigger angle